# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ 9 $K(\pi,1)$-conjecture for some 3D Artin groups
By [ES10], it suffices to show if $\Gamma^{\prime}$ is a complete subgraph of $\Gamma$, then the $K(\pi,1)$-conjecture holds for $A_{\Gamma^{\prime}}$. We claim $\Gamma^{\prime}$ has at most 4 vertices, then the corollary follows by Theorem 9.2, as if the Dynkin diagram $\Lambda$ of $\Gamma$ contain the Dynkin diagram $[3,5,3]$ as an induced subgraph and $W_{\Gamma}$ is a 3-dimensional hyperbolic reflection group, then $\Lambda=[3,5,3]$. It remains to show the claim. Let $P\subset\mathbb{H}^{3}$ be a fundamental domain with respect to the reflection group $W_{\Gamma}$ such that $W_{\Gamma}$ is generated by orthogonal reflection according the codimension 1 faces $\{P_{i}\}_{i=1}^{k}$ of $P$ (some of the vertices of $P$ might be in the boundary at infinity of $\mathbb{H}^{3}$). The support of $P_{i}$ is defined to be the hyperplane in $\mathbb{H}^{3}$ that contains $P_{i}$. We define another graph $\Gamma_{0}$, whose vertices are in 1-1 correspondence with $\{P_{i}\}_{i=1}^{k}$, and two vertices are adjacent if the associated codimension 1 faces intersect in a codimension 2 face. By [And70], two vertices of $\Gamma_{0}$ are adjacent if and only the support of the associated codimension 1 faces have non-empty intersection. Thus $\Gamma_{0}\cong\Gamma$. It follows from the definition of $\Gamma_{0}$ that it is a planar graph. Thus any complete subgraph in $\Gamma$, hence in $\Gamma_{0}$, has $\leq 4$ vertices, as desired. ∎
10 Artin groups with complete bipartite Dynkin diagrams
The goal of this section is to prove $K(\pi,1)$-conjecture for Artin groups with complete bipartite Dynkin diagrams.
@@ 10 Artin groups with complete bipartite Dynkin diagrams
Suppose $\Lambda$ is the Dynkin diagram of type $D_{4}$, and take two leaf vertices $a$ and $b$ from $\Lambda$. The $(a,b)$-subdivision of $\Delta_{\Lambda}$ with its vertex set endowed with the order introduced in Definition 3.30, is a bowtie free and downward flag poset.
@@
Take an embedded 4-cycle $x_{1}y_{1}x_{2}y_{2}$ in $\Delta^{\prime}$ with $t(x_{i})=1$ for $i=1,2$ and $t(y_{i})=4$ for $i=1,2$. Then both $y_{1}$ and $y_{2}$ are of type $\hat{c}$, and each $x_{i}$ is of type $\hat{a}$ or $\hat{b}$. If $x_{1}$ and $x_{2}$ have different types, then it follows from [Hua24, Lemma 4.4] that $x_{1}$ and $x_{2}$ are adjacent in $\Delta_{\Lambda}$. Then there is a vertex of type $m$ in $\Delta^{\prime}$ which is adjacent to each of $\{x_{1},x_{2},y_{1},y_{2}\}$ in $\Delta^{\prime}$, as desired. If $x_{1}$ and $x_{2}$ have same type, then Theorem 3.21 implies there is a vertex of type $\hat{d}$ which is adjacent to each of $\{x_{1},x_{2},y_{1},y_{2}\}$ in $\Delta_{\Lambda}$, hence in $\Delta^{\prime}$, as desired. Thus Lemma 3.24 (3) is verified and the bowtie free part of the lemma follows.
The downward flagness follows from [Hua24, Corollary 7.7]. ∎
@@
Suppose $\Lambda$ is a tripod Dynkin diagram, with a central vertex $a$ and three leave vertices $b_{1},b_{2},b_{3}$. Then for any $i\neq j$, then $(b_{i},b_{j})$-subdivision of $\Delta_{\Lambda}$ is bowtie free and downward flag.
@@
Let $\Lambda$ be a linear Dynkin diagram with vertex set $\{a,b,c\}$. Let $\Delta=\Delta_{\Lambda}$ be the associated Artin complex, with the order on its vertex set induced from $a<b<c$. Then
@@
For Assertion (1), the case of $m_{ab}=4$ and $m_{ab}=5$ follows from Theorem 3.17, Theorem 3.22 and Theorem 7.1. We assume $m_{ab}\geq 6$. The bowtie free condition follows from [Hua23, Corollary 9.13] and [Hua23, Lemma 6.13]. For the downward flagness, by the same argument as in the proof of Theorem 7.1, it suffices to show if we have $\{x_{i}\}_{i=1}^{3}$ of type $\hat{c}$ which are pairwisely lower bounded, then they have a common lower bound. Also we can assume for $i\in\mathbb{Z}/2\mathbb{Z}$, a lower bound of $x_{i}$ and $x_{i+1}$ is $y_{i}$ and $y_{i}$ is of type $\hat{a}$. We metrize triangles in $\Delta_{\Lambda}$ as flat triangles with angle $\pi/6$ at vertices of type $\hat{c}$, angle $\pi/2$ at vertices of type $\hat{b}$ and angle $\pi/3$ at vertices of type $\hat{a}$. By [AS83, Lemma 6], $\Delta_{\Lambda}$ is locally CAT$(0)$, hence CAT$(0)$.
@@
Let $\omega$ be the loop $x_{1}y_{1}x_{2}y_{2}x_{3}y_{3}$ in $\Delta$. We can assume without loss of generality that $\omega$ is an embedded $6$-cycle. Let $\mathbb{D}\to\Delta$ be a minimal area singular disk diagram (Section 2.8) for $\omega$. We endow $\mathbb{D}$ be the induced metric from $\Delta$, and by slightly abusing the notation, we use $x_{i}$ (resp. $y_{i}$) to denote the point in the boundary cycle of $\mathbb{D}$ mapping to $x_{i}$ (resp. $y_{i}$). For $v\in\mathbb{D}^{(0)}$, let $\kappa(v)$ be the quantity defined in Section 2.8. Then $\kappa(y_{i})\leq\pi/3$ for $1\leq i\leq 3$, and $\kappa(v)\leq 0$ for any interior vertex $x\in\mathbb{D}$. Thus (2.15) implies that $\sum_{i=1}^{3}\kappa(x_{i})\geq\pi$.
@@
Note that if $\kappa(x_{i})$ is positive, then the only possible value for $\kappa(x_{i})$ is $\pi/3$ and $32\pi/$ 3. If $\kappa(x_{i})=2\pi/3$, then there is a vertex $x^{\prime}_{i}$ of type $\hat{b}$ such that $x_{i}$ is adjacent to both $y_{i}$ and $y_{i+1}$, and $\overline{y_{i}x^{\prime}_{i}}$ and $x_{i}y_{i+1}$ fit together to form a geodesic in $\Delta$. So if $\kappa(x_{i})=2\pi/3$ for all $i$, then $x^{\prime}_{1}y_{1}x^{\prime}_{2}y_{2}x^{\prime}_{3}y_{3}$ form a geodesic triangle in $\Delta$ with apexes at $\{y_{i}\}_{i=1}^{3}$, which is impossible. If $\kappa(x_{i})=2\pi/3$ for exactly two of $\{x_{i}\}_{i=1}^{3}$, say $x_{1}$ and $x_{2}$, then $x^{\prime}_{1}y_{1}x^{\prime}_{2}y_{2}x_{3}y_{3}$ form a geodesic 4-gon in $\Delta$, with apexes at $y_{1},y_{2},x_{3}$ and $y_{3}$. Now we consider the subdiagram $\mathbb{D}^{\prime}$ of $\mathbb{D}$ for this 4-gon. Then for $\mathbb{D}^{\prime}$, $\kappa(x_{3})\leq\pi/3$, $\kappa(y_{i})\leq 2\pi/3$ for $i=2,3$, $\kappa(x^{\prime}_{i})=0$ for $i=1,2$ (as the link of type $\hat{b}$ vertices in $\Delta$ is complete bipartite), and $\kappa(y_{1})\leq\pi/3$. Thus by (2.15), we must $\kappa(y_{i})=2\pi/3$ for $i=2,3$ and $\kappa(y_{1})=\pi/3$. This implies that $x_{3}$ is adjacent in $\Delta$ to $y_{1}$. Thus $y_{1}$ is a common lower bound for $\{x_{1},x_{2},x_{3}\}$. If $\kappa(x_{i})=2\pi/3$ for exactly one of $\{x_{i}\}_{i=1}^{3}$, say $x_{1}$, then $\sum_{i=1}^{3}\kappa(x_{i})\leq 4\pi/3$, which implies that $\sum_{i=1}^{3}\kappa(y_{i})\geq 2\pi/3$. Thus $\kappa(y_{i})=\pi/3$ for at least two of $\{y_{i}\}_{i=1}^{3}$. Then there is some $y_{i}$ adjacent to $x_{1}$, say $y_{1}$, satisfying $\kappa(y_{1})=\pi/3$. Thus $x_{2}$ is adjacent in $\Delta$ to $y_{3}$, and we finish as before. It remains to consider that $\kappa(x_{i})\leq\pi/3$ for each $i$. Then (2.15) implies that $\kappa(x_{i})=\kappa(y_{i})=\pi/3$ for each $i$ and the disk diagram $\mathbb{D}$ is flat. It follows from the combinatorial structure of flat diagram that there is a vertex $z\in\Delta$ of type $\hat{a}$ such that $z$ is adjacent to $x_{i}$ for $1\leq i\leq 3$, as desired.
@@
Suppose $\Lambda$ is a Dynkin diagram which is a star (i.e. it is a union of edges emanating from the same vertex) with $\geq 4$ vertices. Suppose each tripod subgraph of $\Lambda$ satisfies the conclusion of Proposition 10.2. Suppose the central vertex of $\Lambda$ is $a$. Denote the leaf vertices of $\Lambda$ by $\{b_{i}\}_{i=1}^{n}$ for $n\geq 3$. For $1\leq i,j,k\leq n$, let $\Lambda_{ijk}$ be the subgraph of $\Lambda$ spanned by $a,b_{i},b_{j},b_{k}$. We claim for any pairwise distinct $\{i,j,k\}$, the $(i,j)$-subdivision, $(i,k)$-subdivision and $(j,k)$-subdivision of $\Delta_{\Lambda,\Lambda_{ijk}}$ are all bowtie free and downward flag.
@@
Suppose $\Lambda$ is a Dynkin diagram which is a star (i.e. it is a union of edges emanating from the same vertex) with $\geq 5$ vertices. Suppose each tripod subgraph of $\Lambda$ satisfies the conclusion of Proposition 10.2. Then
@@
Assertion (2) follows from Lemma 10.4. For Assertion (1), let $\mathcal{C}_{2}$ be the class of Dynkin diagrams in Lemma 10.5. Let $\mathcal{C}_{1}=\{\widetilde{D}_{4}\}$. It suffices to verify the first two assumptions of [Hua23, Proposition 7.2]. Assumption 1 of [Hua23, Proposition 7.2] follows from Lemma 10.4 and Proposition 3.34. Assumption 2 is clear. ∎
@@
Let $\Lambda=\Lambda_{m,n}$ be a Dynkin diagram which is a complete bipartite graph that is the join of $m$ vertices and $n$ vertices. Assume $m>1$ and $n>1$. Take a 4-cycle $C$ with its consecutive vertices $\{x_{i}\}_{i=1}^{4}$. Let $f:\Lambda=\Lambda_{m,n}\to C$ be a special folding from $\Lambda_{m,n}$ such that $f^{-1}(x_{i})$ is a single vertex for $i=1,4$, $f^{-1}(x_{2})$ has $n-1$ vertices and $f^{-1}(x_{3})$ has $m-1$ vertices. We view the folded Artin complex $\Delta_{\Lambda,f}$ as a complex of type $S=\{x_{1},x_{2},x_{3},x_{4}\}$, whose vertex set is endowed with a relation induced from the cyclic order $x_{1}<x_{2}<x_{3}<x_{4}<x_{1}$ as explained before Theorem 3.28. Then the link of each vertex of the folded Artin complex $\Delta_{\Lambda,f}$ is bowtie free.
@@
Now we assume at least one of $m$ and $n$, say $n$, is $>2$. Take $v\in\Delta_{\Lambda,f}$ be a vertex of type $\hat{x}_{3}$. Then $\operatorname{lk}(v,\Delta_{\Lambda,f})\cong\Delta_{\Lambda_{3},f}$. Note that $\Lambda_{3}$ is a star. By Lemma 10.5 and [Hua23, Proposition 6.15], $\Delta_{\Lambda_{3}}$ satisfies the labeled four wheel condition. Hence the same holds for $\Delta_{\Lambda_{3},f}$ by Lemma 3.26. Thus $\Delta_{\Lambda_{3},f}$ is bowtie free by Lemma 3.19 and Lemma 3.25. Similarly, we know $\operatorname{lk}(v,\Delta_{\Lambda,f})$ is bowtie free if $v$ is of type $\hat{x}_{2}$.
It remains to consider the case $v$ is of type $\hat{x}_{4}$ or $\hat{x}_{1}$. We will only treat $v$ being of type $\hat{x}_{4}$, as the other case is similar. Note that $\operatorname{lk}(v,\Delta_{\Lambda,f})\cong\Delta_{\Lambda_{4},f}$. Consider a special folding $f^{\prime}:\Lambda_{4}\to C$ such that $(f^{\prime})^{-1}(x_{i})=f^{-1}(x_{i})$ for $i=1,3$, $(f^{\prime})^{-1}(x_{4})$ is one vertex and $(f^{\prime})^{-1}(x_{2})$ has $n-2$ vertices. This is possible as $n>2$. By induction, $\Delta_{\Lambda_{4},f^{\prime}}$ satisfies the conclusion of the claim. Then $\Delta_{\Lambda_{4},f^{\prime}}$, viewed as a complex of type $S=\{x_{1},x_{2},x_{3},x_{4}\}$, satisfies the assumption of Theorem 3.28. Hence Lemma 3.29 applies to $\Delta_{\Lambda_{4},f^{\prime}}$. Now we define an embedding
@@
Suppose $\Lambda$ is a Dynkin diagram which is a complete bipartite graph, and suppose $\Lambda$ has $\geq 5$ vertices if $\Lambda$ is a star. Suppose each tripod subgraph of $\Lambda$ satisfies the conclusion of Proposition 10.2. Then $\Delta_{\Lambda}$ is contractible.

### v3

@@ 10 $K(\pi,1)$-conjecture for some 3D Artin groups
By [ES10], it suffices to show if $\Lambda^{\prime}$ is an induced subgraph of $\Lambda$ which is free-of-infinity, then the $K(\pi,1)$-conjecture holds for $A_{\Lambda^{\prime}}$. We claim $\Lambda^{\prime}$ has at most 4 nodes. The corollary follows from this claim and Theorem 10.3, as if $\Lambda$ contain the Coxeter diagram $[3,5,3]$ as an induced subgraph and $W_{\Lambda}$ is a 3-dimensional hyperbolic reflection group, then $\Lambda=[3,5,3]$. It remains to show the claim. Let $P\subset\mathbb{H}^{3}$ be a fundamental domain with respect to the reflection group $W_{S}$ such that $W_{S}$ is generated by orthogonal reflection according the codimension 1 faces $\{P_{i}\}_{i=1}^{k}$ of $P$ (some of the vertices of $P$ might be in the boundary at infinity of $\mathbb{H}^{3}$). The support of $P_{i}$ is defined to be the hyperplane in $\mathbb{H}^{3}$ that contains $P_{i}$. We define another graph $\Gamma$, whose vertices are in 1-1 correspondence with $\{P_{i}\}_{i=1}^{k}$, and two vertices are adjacent if the associated codimension 1 faces intersect in a codimension 2 face. Then $\Gamma$ is planar. By [And70], two vertices of $\Gamma$ are adjacent if and only the support of the associated codimension 1 faces have non-empty intersection. It follows that nodes of $\Lambda^{\prime}$ span a complete subgraph of $\Gamma$, as $\Lambda^{\prime}$ is free-of-infinity. Thus the claim follows from the planarity of $\Gamma$. ∎
11 Artin groups with complete bipartite Coxeter diagrams
The goal of this section is to prove $K(\pi,1)$-conjecture for Artin groups with complete bipartite Coxeter diagrams. We recall an additional tool from [Hua24a] in Section 11.1 and prove the main theorem in Section 11.2.
11.1 Folded Artin complexes
We recall a variation of Artin complexes and relative Artin complexes introduced in [Hua24a]. Let $\Lambda,\Lambda^{\prime}$ be a Coxeter diagram. A special folding is a surjective graph morphism $f:\Lambda\to\Lambda^{\prime}$ (i.e. it map nodes to nodes and edges to edges) such that if nodes $x$ and $y$ are adjacent in $\Lambda^{\prime}$, then each node in $f^{-1}(x)$ is adjacent in $\Lambda$ to every node in $f^{-1}(y)$.
Given a special folding $f:\Lambda\to\Lambda^{\prime}$, we also use $f:S\to S^{\prime}$ to denote the induced map on the set of nodes. We define the folded Artin complex $\Delta_{\Lambda,f}$ as follows. Vertices of $\Delta_{\Lambda,f}$ are in 1-1 correspondence with left cosets of $A_{S\setminus\{f^{-1}(s^{\prime})\}}$ in $A_{S}$. A collection of vertices span a simplex if the intersection of the associated collection of left cosets is non-empty. By [GP12, Lemma 4.7 and Proposition 4.5], $\Delta_{\Lambda,f}$ is a flag complex.
Lemma 11.1.
[Hua24a, Lemma 10.3] Let $f:\Lambda\to\Lambda^{\prime}$ be a special folding with $\Lambda^{\prime}$ connected. Suppose $\Lambda^{\prime\prime}\subset\Lambda^{\prime}$ is an induced linear subgraph with consecutive nodes $\{s^{\prime}_{i}\}_{i=1}^{n}$. For simplicity we will say a vertex of $Y=\Delta_{\Lambda,f,\Lambda^{\prime\prime}}$ has type $i$ if it has type $\hat{s}^{\prime}_{i}$. We define a relation in the vertex set $V$ of $Y$ as follows. For two vertices $x,y\in Y$, we put $x<y$ if they are adjacent in $Y$ and $\operatorname{Type}(x)<\operatorname{Type}(y)$.
Lemma 11.2.
[Hua24a, Lemma 10.5] Suppose $f:\Lambda\to\Lambda^{\prime}$ be a special folding between two trees. If $\Delta_{\Lambda}$ satisfies the labeled 4-wheel condition, then $\Delta_{\Lambda,f}$ satisfies the labeled 4-wheel condition.
11.2 Proof of contractibility
@@
Suppose $\Lambda$ is the Coxeter diagram of type $D_{4}$, and take two leaf nodes $a$ and $b$ from $\Lambda$. The $(a,b)$-subdivision of $\Delta_{\Lambda}$ with its vertex set endowed with the order introduced in Definition 2.21, is a bowtie free and downward flag poset.
@@
Take an embedded 4-cycle $x_{1}y_{1}x_{2}y_{2}$ in $\Delta^{\prime}$ with $t(x_{i})=1$ for $i=1,2$ and $t(y_{i})=4$ for $i=1,2$. Then both $y_{1}$ and $y_{2}$ are of type $\hat{c}$, and each $x_{i}$ is of type $\hat{a}$ or $\hat{b}$. If $x_{1}$ and $x_{2}$ have different types, then it follows from [Hua24b, Lemma 4.4] that $x_{1}$ and $x_{2}$ are adjacent in $\Delta_{\Lambda}$. Then there is a vertex of type $m$ in $\Delta^{\prime}$ which is adjacent to each of $\{x_{1},x_{2},y_{1},y_{2}\}$ in $\Delta^{\prime}$, as desired. If $x_{1}$ and $x_{2}$ have same type, then Theorem 2.36 implies there is a vertex of type $\hat{d}$ which is adjacent to each of $\{x_{1},x_{2},y_{1},y_{2}\}$ in $\Delta_{\Lambda}$, hence in $\Delta^{\prime}$, as desired. Thus Lemma 8.1 (3) is verified and the bowtie free part of the lemma follows.
The downward flagness follows from [Hua24b, Corollary 7.7]. ∎
@@
Suppose $\Lambda$ is a tripod Coxeter diagram, with a central node $a$ and three leave nodes $b_{1},b_{2},b_{3}$. Then for any $i\neq j$, then $(b_{i},b_{j})$-subdivision of $\Delta_{\Lambda}$ is bowtie free and downward flag.
@@
Let $\Lambda$ be a linear Coxeter diagram with its nodes $\{a,b,c\}$. Let $\Delta=\Delta_{\Lambda}$ be the associated Artin complex, with the order on its vertex set induced from $a<b<c$. Then
@@
For Assertion (1), the case of $m_{ab}=4$ and $m_{ab}=5$ follows from Theorem 2.35, Theorem 2.37 and Theorem 7.1. We assume $m_{ab}\geq 6$. The bowtie free condition follows from [Hua24a, Corollary 9.13] and [Hua24a, Lemma 6.13]. For the downward flagness, by the same argument as in the proof of Theorem 7.1, it suffices to show if we have $\{x_{i}\}_{i=1}^{3}$ of type $\hat{c}$ which are pairwisely lower bounded, then they have a common lower bound. Also we can assume for $i\in\mathbb{Z}/2\mathbb{Z}$, a lower bound of $x_{i}$ and $x_{i+1}$ is $y_{i}$ and $y_{i}$ is of type $\hat{a}$. We metrize triangles in $\Delta_{\Lambda}$ as flat triangles with angle $\pi/6$ at vertices of type $\hat{c}$, angle $\pi/2$ at vertices of type $\hat{b}$ and angle $\pi/3$ at vertices of type $\hat{a}$. By [AS83, Lemma 6], $\Delta_{\Lambda}$ is locally CAT$(0)$, hence CAT$(0)$.
@@
Let $\omega$ be the loop $x_{1}y_{1}x_{2}y_{2}x_{3}y_{3}$ in $\Delta$. We can assume without loss of generality that $\omega$ is an embedded $6$-cycle. Let $\mathbb{D}\to\Delta$ be a minimal area singular disk diagram ([Hua24a, Section 3.6]) for $\omega$. We endow $\mathbb{D}$ be the induced metric from $\Delta$, and by slightly abusing the notation, we use $x_{i}$ (resp. $y_{i}$) to denote the point in the boundary cycle of $\mathbb{D}$ mapping to $x_{i}$ (resp. $y_{i}$). For $v\in\mathbb{D}^{(0)}$, let $\kappa(v)$ be the quantity defined in [Hua24a, Section 3.6]. Then $\kappa(y_{i})\leq\pi/3$ for $1\leq i\leq 3$, and $\kappa(v)\leq 0$ for any interior vertex $x\in\mathbb{D}$. Thus [Hua24a, Equation 3.17] implies that $\sum_{i=1}^{3}\kappa(x_{i})\geq\pi$.
@@
Note that if $\kappa(x_{i})$ is positive, then the only possible value for $\kappa(x_{i})$ is $\pi/3$ and $32\pi/$ 3. If $\kappa(x_{i})=2\pi/3$, then there is a vertex $x^{\prime}_{i}$ of type $\hat{b}$ such that $x_{i}$ is adjacent to both $y_{i}$ and $y_{i+1}$, and $\overline{y_{i}x^{\prime}_{i}}$ and $x_{i}y_{i+1}$ fit together to form a geodesic in $\Delta$. So if $\kappa(x_{i})=2\pi/3$ for all $i$, then $x^{\prime}_{1}y_{1}x^{\prime}_{2}y_{2}x^{\prime}_{3}y_{3}$ form a geodesic triangle in $\Delta$ with apexes at $\{y_{i}\}_{i=1}^{3}$, which is impossible. If $\kappa(x_{i})=2\pi/3$ for exactly two of $\{x_{i}\}_{i=1}^{3}$, say $x_{1}$ and $x_{2}$, then $x^{\prime}_{1}y_{1}x^{\prime}_{2}y_{2}x_{3}y_{3}$ form a geodesic 4-gon in $\Delta$, with apexes at $y_{1},y_{2},x_{3}$ and $y_{3}$. Now we consider the subdiagram $\mathbb{D}^{\prime}$ of $\mathbb{D}$ for this 4-gon. Then for $\mathbb{D}^{\prime}$, $\kappa(x_{3})\leq\pi/3$, $\kappa(y_{i})\leq 2\pi/3$ for $i=2,3$, $\kappa(x^{\prime}_{i})=0$ for $i=1,2$ (as the link of type $\hat{b}$ vertices in $\Delta$ is complete bipartite), and $\kappa(y_{1})\leq\pi/3$. Thus by [Hua24a, Equation 3.17], we must $\kappa(y_{i})=2\pi/3$ for $i=2,3$ and $\kappa(y_{1})=\pi/3$. This implies that $x_{3}$ is adjacent in $\Delta$ to $y_{1}$. Thus $y_{1}$ is a common lower bound for $\{x_{1},x_{2},x_{3}\}$. If $\kappa(x_{i})=2\pi/3$ for exactly one of $\{x_{i}\}_{i=1}^{3}$, say $x_{1}$, then $\sum_{i=1}^{3}\kappa(x_{i})\leq 4\pi/3$, which implies that $\sum_{i=1}^{3}\kappa(y_{i})\geq 2\pi/3$. Thus $\kappa(y_{i})=\pi/3$ for at least two of $\{y_{i}\}_{i=1}^{3}$. Then there is some $y_{i}$ adjacent to $x_{1}$, say $y_{1}$, satisfying $\kappa(y_{1})=\pi/3$. Thus $x_{2}$ is adjacent in $\Delta$ to $y_{3}$, and we finish as before. It remains to consider that $\kappa(x_{i})\leq\pi/3$ for each $i$. Then [Hua24a, Equation 3.17] implies that $\kappa(x_{i})=\kappa(y_{i})=\pi/3$ for each $i$ and the disk diagram $\mathbb{D}$ is flat. It follows from the combinatorial structure of flat diagram that there is a vertex $z\in\Delta$ of type $\hat{a}$ such that $z$ is adjacent to $x_{i}$ for $1\leq i\leq 3$, as desired.
@@
Suppose $\Lambda$ is a Coxeter diagram which is a star (i.e. it is a union of edges emanating from the same node) with $\geq 4$ nodes. Suppose each tripod subgraph of $\Lambda$ satisfies the conclusion of Proposition 11.4. Suppose the central node of $\Lambda$ is $a$. Denote the leaf nodes of $\Lambda$ by $\{b_{i}\}_{i=1}^{n}$ for $n\geq 3$. For $1\leq i,j,k\leq n$, let $\Lambda_{ijk}$ be the subgraph of $\Lambda$ spanned by $a,b_{i},b_{j},b_{k}$. We claim for any pairwise distinct $\{i,j,k\}$, the $(i,j)$-subdivision, $(i,k)$-subdivision and $(j,k)$-subdivision of $\Delta_{\Lambda,\Lambda_{ijk}}$ are all bowtie free and downward flag.
@@
Suppose $\Lambda$ is a Coxeter diagram which is a star (i.e. it is a union of edges emanating from the same node) with $\geq 5$ nodes. Suppose each tripod subgraph of $\Lambda$ satisfies the conclusion of Proposition 11.4. Then
@@
Assertion (2) follows from Lemma 11.6. For Assertion (1), let $\mathcal{C}_{2}$ be the class of Coxeter diagrams in Lemma 11.7. Let $\mathcal{C}_{1}=\{\widetilde{D}_{4}\}$. It suffices to verify the first two assumptions of [Hua24a, Proposition 7.2]. Assumption 1 of [Hua24a, Proposition 7.2] follows from Lemma 11.6 and Proposition 2.26. Assumption 2 is clear. ∎
We need two auxiliary results (Lemmas 11.8 and 11.9) before moving forward.
@@
([Hua24a, Lemma 4.8]) Let $X$ be as in Theorem 2.19, with all the three assumptions there satisfied. Then for any induced 4-cycle in the 1-skeleton of $X$, there is a vertex $x\in X$ such that $x$ is adjacent to each vertex of this 4-cycle.
The following is a small variation of [Hua24a, Lemma 6.13].
Lemma 11.9.
Suppose $\Lambda$ is a linear Coxeter diagram with its consecutive nodes being $S=\{s_{i}\}_{i=1}^{n}$. Let $X$ be a simplicial complex of type $S$. We endow $S$ with a linear order $s_{1}<s_{2}<\cdots<s_{n}$. Assume the vertex set of $X$ equipped with the relation in Definition 2.15 is a poset. Then $X$ satisfies bowtie free condition if and only if it satisfies the labeled 4-wheel condition.
Proof.
Suppose $X$ satisfies bowtie free condition. Take an induced 4-cycle $x_{1}y_{1}x_{2}y_{2}$ in $X$. We can not have $x_{1}<y_{1}<x_{2}$ or $x_{1}>y_{1}>x_{2}$, otherwise $x_{1}$ is adjacent to $x_{2}$ by our assumption, contradicting that we have an induced 4-cycle. Thus $\{x_{1},y_{1},x_{2},y_{2}\}$ forms a bowtie and the labeled 4-wheel condition follows immediately. Now suppose $X$ satisfies the labeled 4-wheel condition. If we have $\{x_{1},y_{1},x_{2},y_{2}\}$ satisfying $x_{i}<y_{j}$ for $1\leq i,j\leq 2$, then these 4 vertices form a 4-cycle in $X$. If $x_{1}$ and $x_{2}$ are comparable, or $y_{1}$ and $y_{2}$ are comparable, then the bowtie free condition clearly holds for $\{x_{1},y_{1},x_{2},y_{2}\}$. Now we assume $x_{1}$ and $x_{2}$ are not comparable, and $y_{1}$ and $y_{2}$ are not comparable. Then the 4-cycle is an induced 4-cycle. Suppose $x_{1}$ has type $\hat{s}_{x_{1}}$ for node $s_{x_{1}}\in\Lambda$. Similarly we define $s_{x_{2}},s_{y_{1}}$ and $s_{y_{2}}$. We assume without loss of generality that the segment $\Lambda^{\prime}$ from $s_{x_{1}}$ to $s_{y_{2}}$ contain all of $\{s_{x_{1}},s_{x_{2}},s_{y_{1}},s_{y_{2}}\}$. Then the labeled 4-wheel condition implies that there is a vertex $z$ adjacent of each vertex of the 4-cycle such that $z$ has type $\hat{s}_{z}$ with node $s_{z}\in\Lambda^{\prime}$. Clearly $x_{1}<z<y_{2}$. Now we show $x_{2}<z$. If this is not true, as $x_{2}$ and $z$ are adjacent, we must have $x_{2}>z$, then $x_{1}<x_{2}$, contradicting to the assumption that $x_{1}$ and $x_{2}$ are not comparable. Similarly, $z<y_{1}$. Thus the bowtie free condition is satisfied. ∎
Lemma 11.10.
Let $\Lambda=\Lambda_{m,n}$ be a Coxeter diagram which is a complete bipartite graph that is the join of $m$ nodes and $n$ nodes. Assume $m>1$ and $n>1$. Take a 4-cycle $C$ with its consecutive vertices $\{x_{i}\}_{i=1}^{4}$. Let $f:\Lambda=\Lambda_{m,n}\to C$ be a special folding from $\Lambda_{m,n}$ such that $f^{-1}(x_{i})$ is a single node for $i=1,4$, $f^{-1}(x_{2})$ has $n-1$ nodes and $f^{-1}(x_{3})$ has $m-1$ nodes. We view the folded Artin complex $\Delta_{\Lambda,f}$ as a complex of type $S=\{x_{1},x_{2},x_{3},x_{4}\}$, whose vertex set is endowed with a relation induced from the cyclic order $x_{1}<x_{2}<x_{3}<x_{4}<x_{1}$ as explained before Theorem 2.19. Then the link of each vertex of the folded Artin complex $\Delta_{\Lambda,f}$ is bowtie free.
@@
Now we assume at least one of $m$ and $n$, say $n$, is $>2$. Take $v\in\Delta_{\Lambda,f}$ be a vertex of type $\hat{x}_{3}$. Then $\operatorname{lk}(v,\Delta_{\Lambda,f})\cong\Delta_{\Lambda_{3},f}$. Note that $\Lambda_{3}$ is a star. By Lemma 11.7 and [Hua24a, Proposition 6.15], $\Delta_{\Lambda_{3}}$ satisfies the labeled four wheel condition. Hence the same holds for $\Delta_{\Lambda_{3},f}$ by Lemma 11.2. Thus $\Delta_{\Lambda_{3},f}$ is bowtie free by Lemma 11.9 and Lemma 11.1. Similarly, we know $\operatorname{lk}(v,\Delta_{\Lambda,f})$ is bowtie free if $v$ is of type $\hat{x}_{2}$.
It remains to consider the case $v$ is of type $\hat{x}_{4}$ or $\hat{x}_{1}$. We will only treat $v$ being of type $\hat{x}_{4}$, as the other case is similar. Note that $\operatorname{lk}(v,\Delta_{\Lambda,f})\cong\Delta_{\Lambda_{4},f}$. Consider a special folding $f^{\prime}:\Lambda_{4}\to C$ such that $(f^{\prime})^{-1}(x_{i})=f^{-1}(x_{i})$ for $i=1,3$, $(f^{\prime})^{-1}(x_{4})$ is one node and $(f^{\prime})^{-1}(x_{2})$ has $n-2$ nodes. This is possible as $n>2$. By induction, $\Delta_{\Lambda_{4},f^{\prime}}$ satisfies the conclusion of the claim. Then $\Delta_{\Lambda_{4},f^{\prime}}$, viewed as a complex of type $S=\{x_{1},x_{2},x_{3},x_{4}\}$, satisfies the assumption of Theorem 2.19. Hence Lemma 11.8 applies to $\Delta_{\Lambda_{4},f^{\prime}}$. Now we define an embedding
@@
Suppose $\Lambda$ is a Coxeter diagram which is a complete bipartite graph, and suppose $\Lambda$ has $\geq 5$ nodes if $\Lambda$ is a star. Suppose each tripod subgraph of $\Lambda$ satisfies the conclusion of Proposition 11.4. Then $\Delta_{\Lambda}$ is contractible.